Prove that a sequence that diverges to +∞ (resp. −∞) is divergent according

to the previous definition.



1
Expert's answer
2021-02-24T12:24:51-0500

Solution:

We say that "\lim _{n \rightarrow \infty} a_{n}=\infty" if for every number M>0 there is an integer N such that "a_{n}>M \quad \text { whenever } \quad n>N" .

Next, we say that "\lim _{n \rightarrow \infty} a_{n}=-\infty" if for every number M<0 there is an integer N such that"a_{n}<M \quad \text{ whenever }\quad n>N" .

According to the above definition, we have the following:

If "\lim _{n \rightarrow \infty} a_{n}" doesn't exist or is infinite we say the sequence diverges. Note that we will say the sequence diverges to "\infty" if "\lim _{n \rightarrow \infty} a_{n}=\infty" or divergent.

Also, "\lim _{n \rightarrow \infty} a_{n}=-\infty" we will say that the sequence diverges to "-\infty" or divergent.